OpenStudy (anonymous):

Find the inverse of the 1 to 1 funcion.

OpenStudy (anonymous):

*function

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x+2}\]

OpenStudy (anonymous):

put y in the place of f(x) then switch the place of x and y then solve for y

OpenStudy (anonymous):

x^(3)-2 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yay!!!!

OpenStudy (anonymous):

remember to replace the y again with f^-1 (x) to say it's the inverse

OpenStudy (anonymous):

so i would just write it as x^(3)-2 or, y=x^(3)-2 ?

OpenStudy (anonymous):

or f(x)=x^(3)-2 ?

OpenStudy (anonymous):

\[f^{-1}(x)=x^3-2\]

OpenStudy (anonymous):

i see, thx!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

:3